# A Note on the Triple Product Property for Finite Groups with Abelian Normal Subgroups of Prime Index

### v1

@@ Abstract.
Three non-empty subsets $S,T,U$ of a finite group $G$ are said to satisfy the triple product property (TPP) if, for elements $s,s^{\prime}\in S$, $t,t^{\prime}\in T$, $u,u^{\prime}\in U$, the equation $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1$ holds if and only if $s=s^{\prime}$, $t=t^{\prime}$, $u=u^{\prime}$. In this case $(S,T,U)$ is called a TPP triple of $G$ and $|S||T||U|$ is called the size of the triple. The triple product ratio of $G$ can be defined as the quantity $\rho(G):=\frac{\beta(G)}{|G|}$, where $\beta(G)$ is the largest size of a TPP triple of $G$, and a special case of this, the subgroup triple product ratio, is the quantity $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|}$, where $\beta_{0}(G)$ is the largest size of a TPP triple of $G$ composed only of subgroups. There is a conjecture that $\rho(G)\leq\frac{4}{3}$ if $G$ contains a cyclic subgroup of index $2$ [4, Conjecture 7.6]. This note proves a version of this conjecture for subgroups by showing that $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$ if $G$ is any group which contains an abelian normal subgroup of prime index $p$.
@@ 3. Technical Results
Let $H\trianglelefteq G$ be abelian, and $G/H=\{gH\mid g\in G\}$ the quotient group of $G$ by $H$, and $(S,T,U)$ a subgroup TPP triple of $G$. Let $\overline{S},\overline{T},\overline{U}\leq G/H$ be the $H$-supports of $S,T,U$ respectively, and $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$.
@@
(2) If $U_{0}=U\cap H\triangleleft G$ and $xH,yH\in\overline{S}\cap\overline{T}$ such that $xH\neq yH$ then $S_{x}^{-1}T_{x}U_{0}$, $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$, where $S_{x}:=S\cap xH$, $T_{x}:=T\cap xH$.
@@
So the assumption $|S||T||U|>|G|=p|H|$ implies that $S_{0}T_{0}U_{0}$ is a subgroup of $H$ of order $|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{3}}\leq|H|$. Furthermore applying part (1) of Lemma 3.3 shows that $\{S_{x}^{-1}T_{x}U_{0}\}_{xH\in G/H}$ and $\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H}$ are collections of cosets of $S_{0}T_{0}U_{0}$ in $H$ having only the trivial coset $S_{0}T_{0}U_{0}$ in common. As $H$ is maximal in $G$ and $U\not\leq H$ it follows that $U_{0}=U\cap H\lhd\langle U,H\rangle=G$. Now applying part (2) of Lemma 3.3 it follows that $\{S_{x}^{-1}T_{x}U_{0}\}$ is a collection of $p$ cosets of $S_{0}T_{0}U_{0}$ in $H$, and similarly that $T_{0}\lhd G$ and $\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H}$ is a collection of $p$ cosets of $S_{0}T_{0}U_{0}$ in $H$, these collections only intersecting trivially in $S_{0}T_{0}U_{0}$. This means that $S_{0}T_{0}U_{0}$ has at least (2p1)(2p-1) cosets in $H$, i.e. $|H:S_{0}T_{0}U_{0}|=\frac{|H|}{|S_{0}||T_{0}||U_{0}|}\geq 2p-1$, which means that $|S||T||U|=p^{3}|S_{0}||T_{0}||U_{0}|\leq p^{3}\cdot\frac{|H|}{(2p-1)}=p^{3}\cdot\frac{1}{p(2p-1)}|G|$$\frac{p^{2}}{2p-1}|G|$.

### v2

@@ Abstract.
Three non-empty subsets $S,T,U$ of a group $G$ are said to satisfy the triple product property (TPP) if, for elements $s,s^{\prime}\in S$, $t,t^{\prime}\in T$, $u,u^{\prime}\in U$, the equation $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1$ holds if and only if $s=s^{\prime}$, $t=t^{\prime}$, $u=u^{\prime}$. In this case $(S,T,U)$ is called a TPP triple of $G$ and $|S||T||U|$ is called the size of the triple. If $G$ is a finite group the triple product ratio of $G$ can be defined as the quantity $\rho(G):=\frac{\beta(G)}{|G|}$, where $\beta(G)$ is the largest size of a TPP triple of $G$, and a special case of this, the subgroup triple product ratio, is the quantity $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|}$, where $\beta_{0}(G)$ is the largest size of a TPP triple of $G$ composed only of subgroups. There is a conjecture that $\rho(G)\leq\frac{4}{3}$ if $G$ contains a cyclic subgroup of index $2$ [4, Conjecture 7.6]. This note proves a version of this conjecture for subgroups by showing that $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$ if $G$ is any finite group which contains an abelian normal subgroup of prime index $p$.
@@ 3. Technical Results
Let $G$ be a group, $H\trianglelefteq G$ an abelian subgroup, $G/H=\{gH\mid g\in G\}$ the quotient group of $G$ by $H$, and $(S,T,U)$ a subgroup TPP triple of $G$. Let $\overline{S},\overline{T},\overline{U}\leq G/H$ be the $H$-supports of $S,T,U$ respectively, and $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$.
@@
(2) If $U_{0}=U\cap H\trianglelefteq G$ and $xH,yH\in\overline{S}\cap\overline{T}$ such that $xH\neq yH$ then $S_{x}^{-1}T_{x}U_{0}$, $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$, where $S_{x}:=S\cap xH$, $T_{x}:=T\cap xH$.
@@
So the assumption $|S||T||U|>|G|=p|H|$ implies that $S_{0}T_{0}U_{0}$ is a subgroup of $H$ of order $|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{3}}\leq|H|$. Furthermore applying part (1) of Lemma 3.3 shows that $\{S_{x}^{-1}T_{x}U_{0}\}_{xH\in G/H}$ and $\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H}$ are collections of cosets of $S_{0}T_{0}U_{0}$ in $H$ having only the trivial coset $S_{0}T_{0}U_{0}$ in common. As $H$ is maximal in $G$ and $U\not\leq H$ it follows that $U_{0}=U\cap H\trianglelefteq\langle U,H\rangle=G$. Now applying part (2) of Lemma 3.3 it follows that $\{S_{x}^{-1}T_{x}U_{0}\}$ is a collection of $p$ cosets of $S_{0}T_{0}U_{0}$ in $H$, and similarly that $T_{0}\trianglelefteq G$ and $\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H}$ is a collection of $p$ cosets of $S_{0}T_{0}U_{0}$ in $H$, these collections only intersecting trivially in $S_{0}T_{0}U_{0}$. This means that $S_{0}T_{0}U_{0}$ has at least (2p1)(2p-1) cosets in $H$, i.e. $|H:S_{0}T_{0}U_{0}|=\frac{|H|}{|S_{0}||T_{0}||U_{0}|}\geq 2p-1$, which means that $|S||T||U|=p^{3}|S_{0}||T_{0}||U_{0}|\leq p^{3}\cdot\frac{|H|}{(2p-1)}=p^{3}\cdot\frac{1}{p(2p-1)}|G|$$\frac{p^{2}}{2p-1}|G|$.